OpenStudy (anonymous):

How long would it take to get $100,000 if I invest $5000 in an account giving 9.7% interst, compounded continuously? Round answer to nearest tenth

OpenStudy (amistre64):

\[A=Pe^{rt}\] \[A/P=e^{rt}\] \[ln(A/P)=rt\] \[\frac{ln(A/P)}{r}=t\]

OpenStudy (anonymous):

13.4 years?

OpenStudy (amistre64):

maybe, id have to actually plug in the numbers :) lets see what the wolf gets (ln(100,000/5,000))/.097 (ln(100/5))/.097 (ln(20))/.097 = ... not 13.4 http://www.wolframalpha.com/input/?i=%28ln%2820%29%29%2F.097